The addition rule for means extends to sums of any number of random variables. Let’s look at a portfolio containing three mutual funds. The monthly returns on Fidelity Magellan Fund, Fidelity Real Estate Fund, and Fidelity Japan Fund for the 36 months ending in December 2000 had approximately these means: W = Magellan monthly return µW = 1.14% X = Real Estate monthly return µX = 0.16% Y = Japan monthly return µY = 1.59% What is the mean monthly return for a portfolio consisting of 50% Magellan, 30% Real Estate, and 20% Japan?
Answer #1

from given data:

mean monthly return for a portfolio =wi*i =1.14*0.5+0.16*0.3+1.59*0.2 =0.936 %